# TKG-Thinker: Towards Dynamic Reasoning over Temporal Knowledge Graphs via Agentic Reinforcement Learning

### v1

@@ Abstract
Temporal knowledge graph question answering (TKGQA) aims to answer time-sensitive questions by leveraging temporal knowledge bases. While Large Language Models (LLMs) demonstrate significant potential in TKGQA, current prompting strategies constrain their efficacy in two primary ways. First, they are prone to reasoning hallucinations under complex temporal constraints. Second, static prompting limits model autonomy and generalization, as it lack optimization through dynamic interaction with temporal knowledge graphs (TKGs) environments. To address these limitations, we propose TKG-Thinker, a novel agent equipped with autonomous planning and adaptive retrieval capabilities for reasoning over TKGs. Specifically, TKG-Thinker performs in-depth temporal reasoning through dynamic multi-turn interactions with TKGs via a dual-training strategy. We first apply Supervised Fine-Tuning (SFT) with chain-of thought data to instill core planning capabilities, followed by a Reinforcement Learning (RL) stage that leverages multi-dimensional rewards to refine reasoning policies under intricate temporal constraints. Experimental results on benchmark datasets with three open-source LLMs show that TKG-Thinker achieves state-of-the-art performance and exhibits strong generalization across complex TKGQA settings.
@@
Zihao Jiang1, Miao Peng3thanks:   Equal contribution., Zhenyan Shan2, Wenjie Xu1, Ben Liu1, Gong Chen1, Ziqi Gao4, Min Peng2 1School of Computer Science, Wuhan University, China 2School of Artificial Intelligence, Wuhan University 3The Hong Kong University of Science and Technology (Guangzhou) 4Tsinghua Shenzhen International Graduate School, Tsinghua University {jiangzihao,bbcavendish,vingerxu,liuben123,chengongcg,pengm}@whu.edu.cn mpeng885@connect.hkust-gz.edu.cn ziqigao@sz.tsinghua.edu.cn
@@ 1 Introduction
Temporal knowledge graphs (TKGs) organize factual knowledge over time and serve as an essential foundation for a wide range of knowledge-driven applications, such as recommendation systems Li et al. (2025c); Chen et al. (2025) and question answering Liu et al. (2025b); Gao et al. (2024). In TKGs, facts are represented as quadruples (subject, relation, object, timestamp). Building upon this representation, temporal knowledge graph question answering (TKGQA) focuses on answering time-sensitive questions by leveraging the knowledge stored in TKGs. For instance, the question “Which team did Luka Dončić play for on 2025-02-03?” can be answered using the quadruple (Luka Dončić, play for, Los Angeles Lakers, 2025-02-03).
@@ 4.1 Supervised Fine-Tuning for Cold Start
4.2 Online RL with Temporal Tools Call
After supervised fine-tuning stage, we further apply the online reinforcement learning with temporal tool calls, which enhances its multi-hop and time-sensitive reasoning capabilities. To enable effective learning in this setting, we address the problem from three key perspectives: action space, reward design, and training objective.
@@ 4.2.1 Action Space
As we can observe the phenomenon that existing retrievers are insensitive to temporal constraints and LLMs tend to overlook temporal requirements and hallucinate during question decomposition Qian et al. (2025); Guo et al. (2025), we identify the critical reasoning characteristics of temporal questions, and we categorize the reasoning paradigms of temporal reasoning over TKGs into specific taxonomies according to timestamps. Crucially, we formalize the agent’s action space as a collection of temporal functional tools. By doing so, we transform temporal reasoning into actionable primitives that directly mirror the TKG structure, moving beyond the limitations of the LLM’s implicit reasoning capabilities. In detail, our action space includes think, plan, temporal search actions, and answer, with the temporal search actions defined as follows:
@@
Based on this action space and treating TKGs as the environment, temporal search tool returns structured feedback enclosed within <observation> and </observation>. Under this formulation, we adopt a ReAct-style Yao et al. (2023); Liu et al. (2025a) interaction protocol, in which the reasoning process proceeds as a a sequence of planning, internal thought generation, temporal search tool calls, and environment feedback. Formally, the interaction trajectory at step $n$ can be represented as:
$\mathcal{H}_{n}=(\tau_{0},p,\tau_{1},a_{1},o_{1},\ldots,\tau_{n},a_{n-1},o_{n-1}),$ (2)
where $\tau$ denotes the agent’s internal thought, $a$ is an action selected from the temporal tool set, with $p$ being the initial planning action, and $o$ is the observation obtained by executing the action $a$ over the TKGs. Based on the historical trajectory $\mathcal{H}_{n}$, the generation process for the next thought $\tau_{n}$ and action $a_{n}$ can be formulated as:
@@ 4.2.2 Reward Design
Outcome Reward evaluates the correctness of final output answer $a_{\mathrm{pred}}$ by comparing it against the ground truth answer $a_{\mathrm{gold}}$ using rule-based criteria exact match (EM):
@@ 4.2.3 Training Objective
With the multi-dimensional reward formulation defined, we formalize the overall training objective of TKG-Thinker framework. To ensure robust policy optimization, we adopt the RLVR paradigm, which can be instantiated through either PPO or GRPO. The policy $\pi_{\theta}$ is optimized by maximizing the objective function $\mathcal{J}(\theta)$, which encourages trajectories with higher-than-average rewards while maintaining stability via importance sampling and Kullback–Leibler (KL) divergence constraints. The overall objective is defined as:
@@ 6 Conclusion
Despite TKG-Thinker’s strong performance achieved in TKGQA, this work remains several limitations. The current reward mechanism relies heavily on binary indicators and rule-based criteria, such as Exact Match (EM) for outcomes and basic format verification. This outcome-based reward lacks a nuanced evaluation of the intermediate reasoning process. Future iterations could incorporate an LLM Judge with detailed rubrics to qualitatively assess the logical and temporal consistency of the think and plan steps, ensuring that the model understands complex temporal constraints rather than just optimizing for a specific output format. Besides, while the model demonstrates effective multi-step reasoning, the relative simplicity of current datasets and benchmarks—particularly their limited reasoning hops—restricts the training of temporal agents capable of long-range planning and inference. Future work should thus explore more complex synthetic multi-hop tasks and open-world settings to foster greater model robustness.
@@ A.4 Baseline Details
we compare TKG-Thinker against three categories of baselines: (1) PLM-based methods, including BERTDevlin et al. (2019), ALBERTLan et al. (2020), and DistilBERTSanh et al. (2019); (2) Embedding-based methods, such as EmbedKGQASaxena et al. (2020), CronKGQASaxena et al. (2021), and MultiQAChen et al. (2023); and (3) LLM-based methods, including Naive RAG Chen et al. (2024a), ReAct RAGYao et al. (2023), ARIChen et al. (2024b), RTQAGong et al. (2025), and TempAgentQianyiHu et al. (2025), TimR4 Qian et al. (2024), and PoK Qian et al. (2025). For consistency with prior work, we adopt the baseline results reported in Qian et al. (2025), Chen et al. (2024b), and QianyiHu et al. (2025) for comparison.
@@ Appendix D Training Dynamics Details
  • Reasoning Expansion (Steps 15-60): The length gradually increases and stabilizes, suggesting that once the format is mastered, TKG-Thinker learns to generate more substantial reasoning chains and necessary tool calls to solve complex temporal queations.

### v2

@@ Abstract
Temporal knowledge graph question answering (TKGQA) aims to answer time-sensitive questions by leveraging temporal knowledge bases. While Large Language Models (LLMs) demonstrate significant potential in TKGQA, current prompting strategies constrain their efficacy in two primary ways. First, they are prone to reasoning hallucinations under complex temporal constraints. Second, static prompting limits model autonomy and generalization, as it lacks optimization through dynamic interaction with temporal knowledge graph (TKG) environments. To address these limitations, we propose TKG-Thinker, a novel agent equipped with autonomous planning and adaptive retrieval capabilities for reasoning over TKGs. Specifically, TKG-Thinker performs in-depth temporal reasoning through dynamic multi-turn interactions with TKGs via a dual-training strategy. We first apply Supervised Fine-Tuning (SFT) with chain of thought data to instill core planning capabilities, followed by a Reinforcement Learning (RL) stage that leverages multi-dimensional rewards to refine reasoning policies under intricate temporal constraints. Experimental results on benchmark datasets with three open-source LLMs show that TKG-Thinker achieves state-of-the-art performance and exhibits strong generalization across complex TKGQA settings.
@@
Zihao Jiang1, Miao Peng2thanks:   Equal contribution., Zhenyan Shan3, Wenjie Xu1, Ben Liu1, Gong Chen1, Ziqi Gao4, Min Peng3 1School of Computer Science, Wuhan University, China 2The Hong Kong University of Science and Technology (Guangzhou) 3School of Artificial Intelligence, Wuhan University 4Tsinghua Shenzhen International Graduate School, Tsinghua University {jiangzihao,bbcavendish,vingerxu,liuben123,chengongcg,pengm}@whu.edu.cn mpeng885@connect.hkust-gz.edu.cn ziqigao@sz.tsinghua.edu.cn
@@ 1 Introduction
Temporal knowledge graphs (TKGs) organize factual knowledge over time and serve as an essential foundation for a wide range of knowledge-driven applications, such as recommendation systems Li et al. (2025c); Chen et al. (2025) and question answering Liu et al. (2025b); Gao et al. (2024). In TKGs, facts are represented as quadruples (subject, relation, object, timestamp). Building upon this representation, temporal knowledge graph question answering (TKGQA) focuses on answering time-sensitive questions by leveraging the knowledge stored in TKGs. For instance, the question “Which team did Luka Dončić play for on 2025-02-03?” can be answered by the quadruple (Luka Dončić, play for, Los Angeles Lakers, 2025-02-03).
@@ 4.1 Supervised Fine-Tuning for Cold Start
4.2 Online RL with Temporal Tool Calls
After the supervised fine-tuning stage, we further apply the online reinforcement learning with temporal tool calls, which enhances its multi-hop and time-sensitive reasoning capabilities. To enable effective learning in this setting, we address the problem from three key perspectives: action space, reward design, and training objective.
@@ 4.2.1 Action Space
As we can observe, existing retrievers are insensitive to temporal constraints and LLMs tend to overlook temporal requirements and hallucinate during question decomposition Qian et al. (2025); Guo et al. (2025), we identify the critical reasoning characteristics of temporal questions, and we categorize the reasoning paradigms of temporal reasoning over TKGs into specific taxonomies according to timestamps. Crucially, we formalize the agent’s action space as a collection of temporal functional tools. By doing so, we transform temporal reasoning into actionable primitives that directly mirror the TKG structure, moving beyond the limitations of the LLM’s implicit reasoning capabilities. In detail, our action space includes think, plan, temporal search actions, and answer, with the temporal search actions defined as follows:
@@
Based on this action space and treating TKGs as the environment, temporal search tool returns structured feedback enclosed within <observation> and </observation>. Under this formulation, we adopt a ReAct-style Yao et al. (2023); Liu et al. (2025a) interaction protocol, in which the reasoning process proceeds as a sequence of planning, internal thought generation, temporal search tool calls, and environment feedback. Formally, the interaction trajectory at step $n$ can be represented as:
$\mathcal{H}_{n}=(\tau_{0},p,\tau_{1},a_{1},o_{1},\ldots,\tau_{n-1},a_{n-1},o_{n-1}),$ (2)
where $\tau$ denotes the agent’s internal thought, $a$ is an action selected from the temporal search actions, with $p$ being the initial planning action, and $o$ is the observation obtained by executing the action $a$ over the TKGs. Based on the historical trajectory $\mathcal{H}_{n}$, the generation process for the next thought $\tau_{n}$ and action $a_{n}$ can be formulated as:
@@ 4.2.2 Reward Design
Outcome Reward evaluates the correctness of the final output answer $a_{\mathrm{pred}}$ by comparing it against the ground truth answer $a_{\mathrm{gold}}$ using a rule-based criteria exact match (EM):
@@ 4.2.3 Training Objective
With the multi-dimensional reward formulation defined, we formalize the overall training objective of TKG-Thinker framework. To ensure robust policy optimization, we adopt the RLVR paradigm, which can be instantiated through either PPO or GRPO.
The policy $\pi_{\theta}$ is optimized by maximizing the objective function $\mathcal{J}(\theta)$, which encourages trajectories with higher-than-average rewards while maintaining stability via importance sampling and Kullback–Leibler (KL) divergence constraints. The overall objective is defined as:
@@ 6 Conclusion
Despite TKG-Thinker’s strong performance achieved in TKGQA, this work has several limitations. The current reward mechanism relies heavily on binary indicators and rule-based criteria, such as Exact Match (EM) for outcomes and basic format verification. This outcome-based reward lacks a nuanced evaluation of the intermediate reasoning process. Future iterations could incorporate an LLM Judge with detailed rubrics to qualitatively assess the logical and temporal consistency of the think and plan steps, ensuring that the model understands complex temporal constraints rather than just optimizing for a specific output format. Besides, while the model demonstrates effective multi-step reasoning, the relative simplicity of current datasets and benchmarks—particularly their limited reasoning hops—restricts the training of temporal agents capable of long-range planning and inference. Future work should thus explore more complex synthetic multi-hop tasks and open-world settings to foster greater model robustness.
@@ A.4 Baseline Details
we compare TKG-Thinker against three categories of baselines: (1) PLM-based methods, including BERTDevlin et al. (2019), ALBERTLan et al. (2020), and DistilBERTSanh et al. (2019); (2) Embedding-based methods, such as EmbedKGQASaxena et al. (2020), CronKGQASaxena et al. (2021), and MultiQAChen et al. (2023); and (3) LLM-based methods, including Naive RAG Chen et al. (2024a), ReAct RAGYao et al. (2023), ARIChen et al. (2024b), RTQAGong et al. (2025), and TempAgentQianyiHu et al. (2025), TimeR4 Qian et al. (2024), and PoK Qian et al. (2025). For consistency with prior work, we adopt the baseline results reported in Qian et al. (2025), Chen et al. (2024b), and QianyiHu et al. (2025) for comparison.
@@ Appendix D Training Dynamics Details
  • Reasoning Expansion (Steps 15-60): The length gradually increases and stabilizes, suggesting that once the format is mastered, TKG-Thinker learns to generate more substantial reasoning chains and necessary tool calls to solve complex temporal questions.